We are experiencing an issue with a hypervisor, affecting the following webservers: pemdublinng041 pemdublinng042 pemdublinng043 pemdublinng044 An engineer is currently investigating Update 8:48: All vms are now back online.
Author Archive | Alan Corrigan
We are experiencing performance issues with connections to this database server mssql792int.cp.blacknight.com An engineer is currently investigating. Update 16:15: A faulty disk was identified and will require a maintenance window at 10pm tonight to complete. The server will be offline for short periods between 22:00 and 23:59pm
Pemdublinng023 will be rebooting at 10:30am today to address a hard disk issue. This should take 5-10 minutes to complete. Update 10:38am: All services back online
The following linux shared hosting servers will be rebooted at 7:30 tomorrow morning to migrate to new hardware. This is expected to take 30 minutes to complete. Pemdublinng003 Pemdublinng004 Pemdublinng005 Pemdublinng006 Update 8:15am : This migration has been postponed